# Teacher Encoder-Student Decoder Denoising Guided Segmentation Network for Anomaly Detection

### v3

@@ Abstract.
Visual anomaly detection is a highly challenging task, often categorized as a one-class classification and segmentation problem. Recent studies have demonstrated that the student-teacher (S-T) framework effectively addresses this challenge. However, most S-T frameworks rely solely on pre-trained teacher networks to guide student networks in learning multi-scale similar features, overlooking the potential of the student networks to enhance learning through multi-scale feature fusion. In this study, we propose a novel model named PFADSeg, which integrates a pre-trained teacher network, a denoising student network with multi-scale feature fusion, and a guided anomaly segmentation network into a unified framework. By adopting a unique teacher-encoder and student-decoder denoising mode, the model improves the student network’s ability to learn from teacher network features. Furthermore, an adaptive feature fusion mechanism is introduced to train a self-supervised segmentation network that synthesizes anomaly masks autonomously, significantly increasing detection performance. Evaluated on the MVTec AD dataset, PFADSeg achieves state-of-the-art results with an image-level AUC of 98.9%, a pixel-level mean precision of 76.4%, and an instance-level mean precision of 78.7%.
Anomaly detection, Student-teacher framework, Multi-scale feature fusion, Self-supervised learning, Anomaly segmentation
1. Introduction
@@
The primary objective of anomaly detection is not only to identify anomalous images but also to precisely segment the anomalous pixel regions. It is critical in various computer vision applications, including industrial defect inspection (Roth et al., 2021; Zavrtanik et al., 2021), medical image analysis (Schlegl et al., 2017), and video surveillance (Liu et al., 2018). However, these tasks commonly face significant challenges, such as the scarcity of anomalous samples and the diversity of anomaly types. This variability makes it difficult to collect sufficient data or labels for the wide range of anomaly categories. Furthermore, the differences between normal and anomalous samples are frequently subtle, often masked by irrelevant noise, which complicates the task of simultaneously filtering out noise and accurately identifying anomalous pixel regions.
@@
Figure 1. The traditional S-T framework (a) and our denoising student-teacher distillation framework (b) in the KD mode.
@@
Given the scarcity of anomaly data and the difficulty of obtaining labels, anomaly detection tasks typically rely on unsupervised learning methods, where models are trained solely on normal data. Recently, the knowledge distillation mode (Salehi et al., 2020), particularly the student-teacher (S-T) framework, has demonstrated its effectiveness in anomaly detection tasks (Bergmann et al., 2020; Deng and Li, 2022; Salehi et al., 2020; Wang et al., 2021; Yamada and Hotta, 2021). In this framework, the teacher network is usually pre-trained on large-scale datasets (e.g., ImageNet (Deng et al., 2009)) and possesses well-established weights, while the student network, which shares the same architecture as the teacher network, is trained under the supervision of the teacher using anomaly detection datasets (Bergmann et al., 2019) containing only normal samples. The goal is for the student network to learn feature representations similar to those of the teacher network. Notably, knowledge distillation can also be applied across multiple levels of a feature pyramid (Salehi et al., 2020; Wang et al., 2021; Zhang et al., 2022), enabling the integration of multi-level feature differences and achieving superior performance. However, several issues remain unresolved. First, the lack of anomalous samples during student network training may lead to insufficient representation of feature differences between the teacher and student networks for anomalies in the S-T framework. Second, the student network in most S-T frameworks shares the same architecture as the teacher network, which may limit the student’s ability to effectively learn features extracted by the teacher, particularly in large-scale anomaly detection datasets where training the student network becomes more complex.
Figure 2. Overview of PFADSeg. The training consists of two stages. In stage (a), the improved student network is trained with pseudo-anomalous images (created by blending anomaly masks with normal images), while the teacher network, with fixed weights, processes normal images. The student learns to match features extracted by the teacher. In stage (b), both networks take pseudo-anomalous images as input, with weights fixed. Their normalized outputs are combined via element-wise multiplication and concatenation to form input for the segmentation network, which is trained using the anomaly masks as ground truth.
@@
To address the aforementioned challenges, we introduce a novel knowledge distillation (KD) approach, the denoising student-teacher distillation mode (Fig. 1), where paired normal and anomalous samples are fed into the teacher and student networks, respectively, with the student network adopting a different architecture from the teacher. Based on this mode, we propose a Parallel Feature Aggregation and Denoising for Anomaly Segmentation (PFADSeg, Fig. 2), which consists of a pre-trained teacher network, an improved denoising student network, and a segmentation network. Synthetic anomaly masks are added to normal images to generate inputs for the student network, while the teacher processes the corresponding unaltered normal images. This setup allows the student network to remove noise and learn features aligned with those of the teacher network, leveraging architectural differences to reduce sensitivity to irrelevant noise. The student network, based on ResNet18 (He et al., 2016), is further enhanced to better fuse features across multiple levels. Finally, a trainable segmentation network is introduced to process multi-level feature differences between the teacher and student networks, with synthetic anomaly masks providing supervision during training.
@@
We evaluated the performance of our model using the unsupervised MVTec AD dataset (Bergmann et al., 2019). Extensive experimental results demonstrate that our approach achieves significant improvements in anomaly detection tasks at the image level, pixel level, and instance level. Additionally, we conducted ablation studies to validate the effectiveness of our proposed model enhancements. Compared to previous work (Zhang et al., 2022), our method introduces several improvements to the segmentation network and the denoising student network:
@@
  • We improve the denoising student network by modifying the residual block structure based on ResNet18 (He et al., 2016). Specifically, we introduce an attention mechanism into the residual connections, enabling dynamic feature fusion and allowing the network to simultaneously focus on both globally distributed large objects and locally distributed small objects.
@@ 2. Related Works
Anomaly detection and localization have been explored from various perspectives, with significant progress made in methods involving image reconstruction, parameter density estimation, and memory-based approaches.
@@
In image reconstruction, researchers have used models such as autoencoders (Bergmann et al., 2018), variational autoencoders (VAE) (Baur et al., 2019), and generative adversarial networks (GANs) (Schlegl et al., 2019, 2017). These models are typically trained on normal data, assuming that anomalous images, unseen during training, will not be effectively reconstructed. The pixel-wise difference between the input image and the reconstructed image is used as an anomaly score. However, these methods face challenges such as overgeneralization, where anomalous regions may still be accurately reconstructed, thus reducing the effectiveness of the anomaly score (Pirnay and Chai, 2022). Recent works have addressed this issue by incorporating techniques such as adversarial training or hybrid models, which combine reconstruction with discriminative learning to improve robustness against anomalies.
From the perspective of parameter density estimation, this method assumes that features extracted from normal data follow a specific distribution, typically a multivariate Gaussian distribution (Defard et al., 2021; Li et al., 2021; Rippel et al., 2021). The model estimates the parameters of this distribution using normal data, and during inference, outliers are identified as anomalies. To overcome the restrictive Gaussian assumption, recent advancements have leveraged normalization flow techniques (Gudovskiy et al., 2022; Yu1 et al., 2021), which transform arbitrary feature distributions into Gaussian distributions. This flexible framework enables the estimation of complex data distributions and has shown significant improvements in detecting various anomalies.
@@
Another line of research focuses on memory-based methods (Roth et al., 2021; Yi and Yoon, 2020), which involve constructing a memory bank of normal data during training. During inference, the model retrieves the most similar items from the memory bank based on a given query and calculates anomaly scores based on their similarity. These methods effectively capture fine-grained patterns in normal data and are particularly useful in scenarios where anomalies are subtle or context-dependent. Extensions of memory-based models, such as those incorporating attention mechanisms or dynamic memory updates, further enhance their applicability to real-world anomaly detection tasks.
@@
Knowledge distillation is a method that leverages a pre-trained teacher network and a trainable student network, and it is highly useful in anomaly detection. The student network is typically trained on a dataset consisting of only normal data, and its anomalous feature representations are expected to differ from those of the teacher network. This divergence provides a basis for distinguishing anomalies. Various strategies have been proposed to enhance the detection of different anomaly types. For example, Bergmann et al. (Bergmann et al., 2020) uses ensemble learning to train multiple student networks, exploiting irregularities in their feature representations to detect anomalies more effectively. Similarly, Salehi et al. (Salehi et al., 2020) and Wang et al. (Wang et al., 2021) introduce methods for aligning multi-level feature representations, allowing the model to capture both low-level and high-level anomalies. These methods demonstrate the versatility and adaptability of knowledge distillation in addressing various challenges in anomaly detection. Additionally, to avoid the scenario where both the teacher and student networks have the same architecture and use the same input data, some works (Deng and Li, 2022; Yamada and Hotta, 2021) design the student network with a decoder architecture. Furthermore, Zhang et al. (Zhang et al., 2022) further distinguishes the representations of anomalies between the teacher and student networks within the S-T framework.
@@
Anomaly Simulation. In the context of one-class classification for anomaly detection (AD), where anomalous data is unavailable for training, pseudo-anomalous data can be simulated for supervised training of AD models. Traditional anomaly simulation methods, such as rotation and cropping (Schlegl et al., 2017), are often ineffective at detecting fine-grained anomaly patterns (Li et al., 2021). Zhang et al. (Zhang et al., 2022) adopted the idea proposed in , using two-dimensional Perlin noise to simulate more realistic anomalous images, and utilized the generated anomaly masks as ground truth for the segmentation network. However, this approach may cause the trained segmentation network to overlook the impact of irrelevant noise in normal images when segmenting real anomalous images. This can lead to misclassifying normal regions as anomalous during inference. The attention on irrelevant noise often shifts focus away from true anomalous regions, thus reducing the effectiveness of anomaly detection, especially in the MVTec AD dataset (Bergmann et al., 2019), where this issue is more pronounced in five texture categories.
3. Method
@@
The proposed PFADSeg architecture consists of three main components: a teacher network pre-trained on the ImageNet (Deng et al., 2009) dataset, an improved denoising student network, and a segmentation network. The training process utilizes synthetic anomaly images generated based on the algorithm in (Zavrtanik et al., 2021). Specifically, random 2D Perlin noise is binarized to create an anomaly mask $k$. The masked regions are replaced with a linear combination of the normal image $M_{n}$ and an external data source $A$, producing the synthetic anomaly image $M_{a}$. The transparency factor $\beta$ is sampled from $[0.15,1]$, and $\odot$ represents element-wise multiplication. The process is defined as follows:
(1) $M_{a}=\beta(K\odot A)+(1-\beta)(K\odot M_{n})+(1-K)\odot M_{n}$
@@
Figure 3. Comparison between the original Residual Block (a) in ResNet18 and the proposed PA_Residual Block (b).
Figure 4. Overall architecture of the PCAR module. The feature map is represented by its dimensions $H$, $W$, and $C$, denoting height, width, and the number of channels, respectively. The Softmax operation recalibrates channel attention weights to capture long-range channel dependencies. The symbol $\otimes$ indicates element-wise multiplication, and channel_sum refers to summation along the channel dimension.
3.1. Teacher Encoder-Student Decoder Denoising Network
Our proposed method introduces a modified denoising student network that is fundamentally different from the teacher network. Specifically, both the teacher and student networks utilize paired normal and anomalous images as inputs, but the denoising student network adopts a different architecture from the teacher network. The teacher network we use is a ResNet18 (He et al., 2016) pre-trained on the ImageNet dataset, but only the first three convolutional blocks (i.e., conv1_x, conv2_x, conv3_x) are employed, denoted as $T_{1}$, $T_{2}$, and $T_{3}$, respectively. For the denoising student network, the encoder part is based on an improved ResNet18, which we refer to as PA_ResNet18, where the residual blocks have been enhanced. This version of the student network uses four convolutional blocks, named $PA\_SE^{1}$, $PA\_SE^{2}$, $PA\_SE^{3}$, and $PA\_SE^{4}$. For the decoder part, we employ a reversed PA_ResNet18 (where all downsampling layers are replaced with bilinear upsampling). This decoder also utilizes four convolutional blocks, named $PA\_SD^{1}$, $PA\_SD^{2}$, $PA\_SD^{3}$, and $PA\_SD^{4}$.
@@ 3.1. Teacher Encoder-Student Decoder Denoising Network
Feature fusion is a method of combining features from different layers or branches, often achieved through simple operations like summation or concatenation. However, these simple operations may not always be the best choice. To address this, we draw inspiration from the Attentional Feature Fusion (AFF) module proposed in (Dai et al., 2021). Since anomaly detection is not merely a simple binary classification problem but more importantly involves anomaly region segmentation, we aim to handle feature fusion across different scenarios. We also introduce attention mechanisms and multi-scale context aggregation, allowing the model to aggregate contextual information from various receptive fields. This enables the network to adapt to objects of different scales, focusing on both large objects with global distribution and small objects with local distribution. This is particularly important for anomaly detection, especially when segmenting small anomalous regions.
The improved residual block in our PA_ResNet18 is shown in Fig. 3. We replace the simple summation-based feature fusion in the residual connection of the original Residual Block with the feature fusion method involving the AFF. Rather than simply summing the features, we employ AFF to dynamically fuse them. Additionally, the second $3\times 3$ convolution layer in the original Residual Block is replaced by the Parallel Convolutional Attention Recalibration (PCAR) module we proposed, which will be elaborated in detail in Sec. 3.2. This modification enhances the model’s ability to focus on relevant features and improves its performance in detecting and segmenting anomalies.
We minimize the cosine distance between the features $T_{i}$ and $PA\_SD^{i}$ (where $i=1,2,3$). Let $F_{T_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ be the features extracted by the teacher network, and $F_{S_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ be the features extracted by the decoder part of the student network. The cosine distance is calculated using the (2) and (3), where $j$ and $k$ represent the spatial coordinates on the feature map, with $j=1\ldots H_{i}$ and $k=1\ldots W_{i}$. The total loss is calculated as the sum of the losses for three different feature pairs, as shown in (4). Where, $X_{i}(j,k)$ is the cosine similarity between the features from the teacher and student network; $D_{i}(j,k)$ is the cosine distance, which is used to calculate the dissimilarity between the teacher and student features; $L_{\cos}$ represents the total loss over the three feature pairs, used to guide the training process.
(2) $X_{i}(j,k)=\frac{F_{T_{i}}(j,k)\odot F_{S_{i}}(j,k)}{\left\|F_{T_{i}}(j,k)\right\|_{2}\left\|F_{S_{i}}(j,k)\right\|_{2}}$
(3) $D_{i}(j,k)=1-\sum_{c=1}^{G_{i}}X_{i}(j,k)_{c}$
(4) $L_{\cos}=\sum_{i=1}^{3}\left(\frac{1}{H_{i}W_{i}}\sum_{j,k=1}^{H_{i},W_{i}}D_{i}(j,k)\right)$
3.2. Parallel Convolutional Attention Recalibration Module

### v4

@@ Abstract
Visual anomaly detection is a highly challenging task, often categorized as a one-class classification and segmentation problem. Recent studies have demonstrated that the student-teacher (S-T) framework effectively addresses this challenge. However, most S-T frameworks rely solely on pre-trained teacher networks to guide student networks in learning multi-scale similar features, overlooking the potential of the student networks to enhance learning through multi-scale feature fusion. In this study, we propose a novel model named PFADSeg, which integrates a pre-trained teacher network, a denoising student network with multi-scale feature fusion, and a guided anomaly segmentation network into a unified framework. By adopting a unique teacher-encoder and student-decoder denoising mode, the model improves the student network’s ability to learn from teacher network features. Furthermore, an adaptive feature fusion mechanism is introduced to train a self-supervised segmentation network that synthesizes anomaly masks autonomously, significantly increasing detection performance. Rigorous evaluations on the widely-used MVTec AD dataset demonstrate that PFADSeg exhibits excellent performance, achieving an image-level AUC of 98.9%, a pixel-level mean precision of 76.4%, and an instance-level mean precision of 78.7%.
1 Introduction
@@
The primary objective of anomaly detection is not only to identify anomalous images but also to precisely segment the anomalous pixel regions. It is critical in various computer vision applications, including industrial defect inspection [13, 23], medical image analysis [17], and video surveillance [10]. However, these tasks commonly face significant challenges, such as the scarcity of anomalous samples and the diversity of anomaly types. This variability makes it difficult to collect sufficient data or labels for the wide range of anomaly categories. Furthermore, the differences between normal and anomalous samples are frequently subtle, often masked by irrelevant noise, which complicates the task of simultaneously filtering out noise and accurately identifying anomalous pixel regions.
@@
Figure 1: The traditional S-T framework (a) and our denoising student-teacher distillation framework (b) in the KD mode.
@@
Due to the scarcity of anomaly data and labeling difficulties, anomaly detection often employs unsupervised methods, training models exclusively on normal data. Recently, the knowledge distillation mode [15], particularly the student-teacher (S-T) framework, has demonstrated its effectiveness in anomaly detection tasks [3, 7, 15, 18, 20]. In this framework, a teacher network, typically pre-trained on large datasets (e.g., ImageNet [8]) with well-established weights, guides a student network. The student, often sharing the teacher’s architecture, is trained on anomaly detection datasets [2] containing only normal samples, aiming to mimic the teacher’s feature representations for these samples. Furthermore, knowledge distillation across multiple feature pyramid levels [15, 18, 25] allows for the integration of multi-level feature differences, often leading to enhanced performance. Despite these advancements, several challenges persist. First, the lack of anomalous samples during student network training may lead to insufficient representation of feature differences between the teacher and student networks for anomalies in the S-T framework. Second, the student network in most S-T frameworks shares the same architecture as the teacher network, which may limit the student’s ability to effectively learn features extracted by the teacher, particularly in large-scale anomaly detection datasets where training the student network becomes more complex.
@@
To address the aforementioned challenges, we introduce a novel knowledge distillation (KD) approach, the denoising student-teacher distillation mode (see Fig. 1), where paired normal and anomalous samples are fed into the teacher and student networks, respectively, with the student network adopting a different architecture from the teacher. Based on this mode, we propose a Parallel Feature Aggregation and Denoising for Anomaly Segmentation (PFADSeg, Fig. 2), which consists of a pre-trained teacher network, an improved denoising student network, and a segmentation network. Synthetic anomaly masks are added to normal images to generate inputs for the student network, while the teacher processes the corresponding unaltered normal images. This setup allows the student network to remove noise and learn features aligned with those of the teacher network, leveraging architectural differences to reduce sensitivity to irrelevant noise. The student network, based on ResNet18 [9], is further enhanced to better fuse features across multiple levels. Finally, a trainable segmentation network is introduced to process multi-level feature differences between the teacher and student networks, with synthetic anomaly masks providing supervision during training.
Figure 2: Overview of PFADSeg. The training consists of two stages. In stage (a), the improved student network is trained with pseudo-anomalous images (created by blending anomaly masks with normal images), while the teacher network, with fixed weights, processes normal images. The student learns to match features extracted by the teacher. In stage (b), both networks take pseudo-anomalous images as input, with weights fixed. Their normalized outputs are combined via element-wise multiplication and concatenation to form input for the segmentation network, which is trained using the anomaly masks as ground truth.
@@
We evaluated the performance of our model using the unsupervised MVTec AD dataset [2]. Extensive experimental results demonstrate that our approach achieves significant improvements in anomaly detection tasks at the image level, pixel level, and instance level. Additionally, we conducted ablation studies to validate the effectiveness of our proposed model enhancements. Compared to previous work [25], our method introduces several improvements to the segmentation network and the denoising student network:
@@
  • We improve the denoising student network by modifying the residual block structure based on ResNet18 [9]. Specifically, we introduce an attention mechanism into the residual connections, enabling dynamic feature fusion and allowing the network to simultaneously focus on both globally distributed large objects and locally distributed small objects.
@@ 2 Related Works
Anomaly detection and localization have been explored from various perspectives, with significant progress in methods based on image reconstruction, memory banks, and knowledge distillation.
@@
In image reconstruction, common models include autoencoders [4], variational autoencoders (VAEs) [1], and generative adversarial networks (GANs) [16, 17]. Trained on normal data, these models aim to poorly reconstruct unseen anomalous images. The pixel-wise difference between the input image and the reconstructed image is used as an anomaly score. However, a key challenge is overgeneralization, where anomalous regions might be accurately reconstructed, diminishing the anomaly score’s effectiveness [12]. To mitigate this, recent works incorporate adversarial training or hybrid models that combine reconstruction with discriminative learning for improved anomaly robustness.
@@
Memory-based methods [13, 21] construct a memory bank from normal training data. During inference, anomaly scores are derived from the similarity between a query and its closest matches in the memory bank. These methods effectively capture fine-grained patterns in normal data, especially for subtle or context-dependent anomalies. Extensions incorporating attention mechanisms or dynamic memory updates further enhance their real-world applicability.
@@
Knowledge distillation, leveraging a pre-trained teacher network and a trainable student network, is highly effective for anomaly detection. The student network is typically trained on a dataset consisting of only normal data, and its anomalous feature representations are expected to differ from those of the teacher network. This divergence provides a basis for distinguishing anomalies. Various strategies have been proposed to enhance the detection of different anomaly types. For instance, Bergmann et al. [3] employed ensemble learning with multiple students, leveraging irregularities in their feature representations for more effective anomaly detection. Salehi et al. [15] and Wang et al. [18] introduced multi-level feature alignment to capture diverse anomalies. These methods demonstrate the versatility and adaptability of knowledge distillation in addressing various challenges in anomaly detection. To prevent identical architectures and inputs for teacher and student networks, some works [7, 20] utilize a student-decoder architecture. Furthermore, Zhang et al. [25] further distinguishes the representations of anomalies between the teacher and student networks within the S-T framework.
@@
Anomaly Simulation. For one-class anomaly detection (AD) where training anomalies are absent, simulating pseudo-anomalous data enables supervised model training. Traditional simulation methods, like rotation and cropping [17], often prove ineffective for fine-grained anomaly patterns. Zhang et al. [25] adopted the idea proposed in [23] , using two-dimensional Perlin noise to simulate more realistic anomalous images, and utilized the generated anomaly masks as ground truth for the segmentation network. However, such an approach risks the segmentation network overlooking the impact of irrelevant noise in normal images when segmenting actual anomalies. This can lead to misclassifying normal regions as anomalous during inference. Focusing on irrelevant noise can detract from true anomalies, reducing detection effectiveness, particularly for certain texture categories in datasets like MVTec AD [2].
3 Method
@@
The proposed PFADSeg architecture consists of three main components: a teacher network pre-trained on the ImageNet [8] dataset, an improved denoising student network, and a segmentation network. The training process utilizes synthetic anomaly images generated based on the algorithm in [23]. Specifically, random 2D Perlin noise is binarized to create an anomaly mask $k$. The masked regions are replaced with a linear combination of the normal image $M_{n}$ and an external data source $A$, producing the synthetic anomaly image $M_{a}$. The transparency factor $\beta$ is sampled from $[0.15,1]$, and $\odot$ represents element-wise multiplication. The process is defined as follows:
$M_{a}=\beta(K\odot A)+(1-\beta)(K\odot M_{n})+(1-K)\odot M_{n}$ (1)
@@
3.1 Teacher Encoder-Student Decoder Denoising Network
A key aspect of our approach is the distinct architecture employed by the denoising student network compared to the teacher. The teacher network is based on a ResNet18 [9] pre-trained on ImageNet, utilizing only its first three convolutional blocks (conv1_x, conv2_x, conv3_x), denoted $T_{1}$, $T_{2}$, and $T_{3}$, respectively. In contrast, the denoising student network features an encoder based on an enhanced ResNet18 architecture (termed PA_ResNet18), featuring modified residual blocks. This encoder comprises four convolutional blocks ($PA\_SE^{1}$ to $PA\_SE^{4}$). The student’s decoder is constructed similarly but in reverse (replacing downsampling with bilinear upsampling) and also consists of four convolutional blocks ($PA\_SD^{1}$ to $PA\_SD^{4}$).
@@ 3.1 Teacher Encoder-Student Decoder Denoising Network
Feature fusion, the combination of features from different layers or branches, commonly uses simple operations like summation or concatenation. However, such simple operations can be suboptimal. Seeking more effective fusion, we draw inspiration from the Attentional Feature Fusion (AFF) module [6]. Recognizing that anomaly detection requires precise region segmentation, not just binary classification, our goal is to develop a feature fusion mechanism suitable for diverse anomaly scenarios. Our approach incorporates attention mechanisms and multi-scale context aggregation to effectively gather contextual information across different receptive fields. This enables the network to adapt to objects of different scales, focusing on both large objects with global distribution and small objects with local distribution. Such adaptability is crucial for anomaly detection, particularly for segmenting small anomalous regions accurately.
Figure 3: Comparison between the original Residual Block (a) in ResNet18 and the proposed PA_Residual Block (b).
The architecture of the enhanced residual block used in our PA_ResNet18 is shown in Fig. 3. Instead of the standard summation in the residual connection, we employ the Attentional Feature Fusion (AFF) mechanism for dynamic feature fusion. Furthermore, the second $3\times 3$ convolution layer is replaced by our proposed Parallel Convolutional Attention Recalibration (PCAR) module (detailed in Sec. 3.2). These modifications enhance the block’s ability to focus on relevant features, thereby improving overall anomaly detection and segmentation performance.
The training objective is to minimize the cosine distance between corresponding feature maps from the teacher blocks ($T_{i}$) and the student decoder blocks ($PA\_SD^{i}$) for $i=1,2,3$. Let $F_{T_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ and $F_{S_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ denote these feature maps, respectively. We first compute the cosine similarity $X_{i}(j,k)$ and subsequently the cosine distance $D_{i}(j,k)$ between the feature vectors at each spatial location $(j,k)$ (where $1\leq j\leq H_{i},1\leq k\leq W_{i}$), following Equations (2) and (3). The overall cosine distance loss, $L_{\text{cos}}$, guiding the training, is the sum of these distances across all locations and all three feature pairs ($i=1,2,3$), as formulated in Equation (4).
$X_{i}(j,k)=\frac{F_{T_{i}}(j,k)\odot F_{S_{i}}(j,k)}{\left\|F_{T_{i}}(j,k)\right\|_{2}\left\|F_{S_{i}}(j,k)\right\|_{2}}$ (2)
$D_{i}(j,k)=1-\sum_{c=1}^{C_{i}}X_{i}(j,k)_{c}$ (3)
$L_{\cos}=\sum_{i=1}^{3}\left(\frac{1}{H_{i}W_{i}}\sum_{j,k=1}^{H_{i},W_{i}}D_{i}(j,k)\right)$ (4)
3.2 Parallel Convolutional Attention Recalibration Module